OpenStudy (anonymous):

Find the coordinates of the other endpoint when you are given the midpoint and one of the endpoints. Answer in terms of a decimal rounded to the nearest tenth or an improper fraction in simplest form. P1 M P2 (3, 5) (-2, 0) ( , ) helpppp

OpenStudy (anonymous):

-2=(x-3)/2 0=(y-5)/2 This?

OpenStudy (anonymous):

please show me step by step :/ idk how to do this

OpenStudy (anonymous):

well midpoint=(xf-xi)/2

OpenStudy (anonymous):

Plug and solve for unknown

OpenStudy (anonymous):

@timo , isn't it supposed to be -2=(x+3)/2 ?

OpenStudy (anonymous):

hmm i'll check

OpenStudy (anonymous):

but idk how to solve that -2=(x+3)/2 i need help..

OpenStudy (anonymous):

ah yes it is plus

OpenStudy (anonymous):

midpoint is (averate of the xs, average of the ys)

OpenStudy (anonymous):

just multipy by 2 first and subtract 3

OpenStudy (anonymous):

so it would b 3?

OpenStudy (anonymous):

i get -7 for x

OpenStudy (anonymous):

how??

OpenStudy (anonymous):

-2=(x+3)/2 multiply by 2 both sides -4=x+3 subtract by 3 both sides -7=x :)

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhhhhhhhhh

OpenStudy (anonymous):

duh -.- THANKS!

OpenStudy (anonymous):

for Y it is the same way 0=(y+5)/2 0=y+5 -5=y :)

OpenStudy (anonymous):

LOL :) Who are you and what did you do with seashell :P?

OpenStudy (anonymous):

hahhaaha...

OpenStudy (anonymous):

blonde lazy morning -.-

OpenStudy (anonymous):

but woudnt you multiply the zero by 2?

OpenStudy (anonymous):

Yes and it will make it 0

OpenStudy (anonymous):

Id slap myself

OpenStudy (anonymous):

::P No way I seen you here before :) maybe you do your better work at night :)

OpenStudy (anonymous):

-.- lol